# Stellar Mass Segregation in Dark Matter Halos

### v1

@@ 3.2 Initial conditions
$\rho_{\star}(r)=M_{\star}/(8\pi r_{\star}^{3})~{}\exp(-r/r_{\star})~{},$ (7)
@@
$\rho_{\mathrm{sub}}(r)=M_{\mathrm{sub}}/(2\pi r_{\mathrm{sub}}^{3})~{}(r/r_{\mathrm{sub}})^{-1}(1+r/r_{\mathrm{sub}})^{-3}~{}.$ (8)
@@
$N$-body realizations. We generate equilibrium $N$-body realizations of Eq. 7 in the combined potential of the dark matter and stellar components using using the Eddington-inversion code nbopy (Errani & Peñarrubia, 2020), available online [^5] https://github.com/rerrani/nbopy. We assume that both stellar components have isotropic velocities in the initial conditions. The dark matter subhalo (Eq. 8) is modelled as an analytical background potential. To reduce the impact of Poisson noise on our analysis, for each choice of dynamical-to-stellar mass ratio $\Upsilon_{\mathrm{dyn}}$, we generate $200$ realizations of our UMa3/U1 model, $40$ for Del1, and $10$ for Eri3.
@@ 3.3 $N$-body code
For the UMa3/U1 models, we update the particle tree responsible for the long-range forces with a step size of $\Delta t_{\mathrm{L}}=2^{-10}r_{\mathrm{sub}}^{3/2}(GM_{\mathrm{sub}})^{-1/2}$. With this choice, the period of a circular orbit at the initial half-light radius $r_{\mathrm{h}}/r_{\mathrm{sub}}=1/500$ of a particle that is only subject to long-range forces (including the force due to the analytical dark matter potential) is resolved by ${\approx}270$ steps for the models with $\Upsilon_{\mathrm{dyn}}=10$. To test for numerical convergence, we have decreased and increased this value by factors of 4, with no qualitative impact on our results. Following the recommendation in Wang et al. (2020a, see their Eq. 12 and 41), we choose as reference radii for the separation of short- and long-range forces the values $r_{\mathrm{in,ref}}=\Delta t_{\mathrm{L}}\sigma_{1D}$ and $r_{\mathrm{out,ref}}=10\,r_{\mathrm{in,ref}}$, where by $\sigma_{1D}$ we denote the $N$-body system’s velocity dispersion. Slow-down regularization through sdar is applied to particles below a radius $r_{\mathrm{bin}}=0.8\,r_{\mathrm{in,ref}}$ (the default value in petar). To test for convergence, we have decreased this radius by a factor of 8, again with no qualitative impact on our results.
For the Del1 models, we use the same tree time step and reference radii as for UMa3/U1. Instead for Eri3, we use a shorter tree time step of $\Delta t_{\mathrm{L}}=2^{-12}r_{\mathrm{sub}}^{3/2}(GM_{\mathrm{sub}})^{-1/2}$, which results in better performance at that $N$-body particle number by reducing the number of particles within $r_{\mathrm{in,ref}}$.
Figure 2: Stellar mass segregation in $N$-body realizations of the example systems UMa3/U1 (top), Del1 (center) and Eri1 (bottom). Each system is embedded in a cuspy dark matter halo, with an initial dynamical-to-stellar mass ratio of $\Upsilon_{\mathrm{dyn}}=10$. Shown are $\{x,y\}$ projections of snapshots at $t/\mathrm{Gyr}=0$, $2.5$, $5$, $7.5$ and $10$. The axes are expressed in units of the initial half-light radius $r_{\mathrm{h0}}$. Individual high-mass stars are shown as dark grey points, while low-mass stars are shown in blue. The median half-light radii of high-mass and low-mass stars (computed over a sample of $N$-body realizations, see text) are shown as black and blue circles, respectively. Over a period of $10\,\mathrm{Gyr}$, the half-light radius of the population of low-mass stars expands, while the half-light radius of the high-mass population contracts. Consistent with the relaxation time estimates of Fig. 1, the effect is largest for the UMa3/U1 model, and smallest for the Eri3 model. Dynamically-formed binaries consisting of two high-mass stars are highlighted in red (see Sec. 4.2). A video version of this figure is available as an arXiv ancillary file.
@@
4 Simulations results
We can now turn our attention to the results of our $N$-body experiments. In Sec. 4.1, we describe the simulations of UMa3/U1, Del1 and Eri3 assuming that the stellar component is embedded in a static dark matter halo. Then, in Sec. 4.4, we model the effect of tides on the system by allowing the underlying dark matter potential to evolve with time.
@@ 4.1 Simulations in a static dark matter halo
Figure 2 shows simulation snapshots of the UMa3/U1 (top), Del1 (center) and Eri3 (bottom) models, evolved for $10\,\mathrm{Gyr}$ in a static dark matter halo. These models have an initial dynamical-to-stellar mass ratio of $\Upsilon_{\mathrm{dyn}}=10$ within the (initial) half-light radius. Individual low-mass (high-mass) stars are shown as light-blue (grey) points, respectively. As time progresses, the population of low-mass stars expands, while the population of high-mass stars contracts: Even though the system is highly dark matter-dominated, the stellar population undergoes mass segregation [^6] Mass segregation occurs as close encounters between stars provide a means for the exchange of energy. The system’s tendency towards equipartition of energy results in low-mass stars to preferentially move to less-bound orbits (the low-mass population expands), whereas the high-mass stars move towards more tightly bound orbits (the high-mass population contracts), see e.g. Spitzer (1987, chapter 1.3). driven by collisional relaxation. Blue and black circles in Figure 2 show the median half-light radii of the low-mass (high-mass) population, with the median computed from the sample of all $N$-body realizations. Mass segregation progresses fastest for UMa3/U1 and slowest for Eri3, consistent with the relaxation times estimated in Fig. 1.
The detailed time evolution of the models with $\Upsilon_{\mathrm{dyn}}=10$ is shown in Fig. 3. The top panels show the evolution of the (median) half-light radii of the populations of low-mass (blue) and high-mass (black) stars. For the case of UMa3/U1, after $10\,\mathrm{Gyr}$ of evolution, the half-light radius of the population of low-mass stars is twice as large as the half-light radius of high-mass stars. For Del1, the difference reduces to ${\sim}40$ per cent, and for Eri3 to ${\sim}10$ per cent. The evolution of individual $N$-body realizations is shown in lighter shades in the background: For systems akin to UMa3/U1, Poisson noise driven by the low number of stars in the system substantially complicates a detection of mass segregation. The bottom panels of Fig. 3 show the evolution of the (3D) stellar velocity dispersion $\langle v^{2}\rangle=\sum v^{2}/N_{\star}$. As the half-light radius of the population of low-mass stars expands, the velocity dispersion heats up. Vice-versa, as the population of high-mass stars contracts, the population cools down.
@@ 4.2 Dynamical formation and disruption of binaries
As the population of high-mass stars contracts and cools, we note the dynamical formation of stellar binary systems in our simulations. We identify binaries as pairs of stars with Keplerian binding energy $E_{\mathrm{bin}}<0$ and an orbital period $T_{\mathrm{bin}}$ that is shorter than the (circular) period $T_{\mathrm{com}}$ of the pair’s center of mass within the dark matter subhalo. Expressed in terms of densities, this condition implies that the mean stellar density within the semi-major axis $a_{\mathrm{bin}}$ of a binary exceeds the mean density of the dark halo within a radius equal to that of the binary’s center of mass , i.e $(m_{i}+m_{j})/a_{\mathrm{bin}}^{3}>M_{\mathrm{sub}}({<}r_{\mathrm{com}})/r_{\mathrm{com}}^{3}$, where by $m_{i}$ and $m_{j}$ we denote the masses of the two binary components. We count the number of binaries at each snapshot in the simulation. For better statistics, we stack $2000$ realizations of the UMa3/U1 model with an initial dynamical-to-stellar mass ratio of $\Upsilon_{\mathrm{dyn}}=10$. Note that our initial conditions are created by drawing individual stars from the underlying distribution function; any binaries present in the initial conditions just arise from this random sampling.
@@
Figure 4 shows the binary fraction $f_{\mathrm{bin}}\equiv N_{\mathrm{bin},i}/N_{i}$, defined here as the number $N_{\mathrm{bin},i}$ of high-mass stars that are in binary systems, normalized by the total number of high-mass stars $N_{i}$. A grey band show the binary fraction considering only pairs of two high-mass stars, whereas the blue band corresponds to binaries that consist of a high-mass star and a low-mass star. As the population of high-mass stars contracts and cools due to mass segregation, the number of dynamically formed massive–massive binaries increases. At the same time, as the population of low-mass stars expands and heats up, the number of massive–low mass binaries drops.
@@
Some intuition in the processes driving the formation and disruption of dynamical binaries can be gained from the statistical theory of gravitational capture, developed to estimate the number of gravitationally trapped (mass-less) tracer particles around a point-mass perturber orbiting in a smooth dark matter halo (Peñarrubia, 2023). In our collisional $N$-body simulations, all stars are massive particles, and complicated three- or multi-body interaction between stars and the halo likely contribute to the formation and disruption processes. Nevertheless, as we will show in the following, the statistical theory of gravitational capture provides accurate estimates for the binary fractions found in our simulations. Building upon equation 4 in Peñarrubia (2021), we estimate the binary fraction through [^7] We obtain our equation 9 from equation 4 in Peñarrubia (2021) by approximating the mean number density of field stars trough $\bar{n}_{j}=3N_{j}/(8\pi r_{\mathrm{h}j}^{3})$. We then substitute the mass of the perturber by the mass of the binary system, $m_{i}+m_{j}$. Finally, we compute the number of bound field stars within a radius $a_{\mathrm{max}}=r_{\mathrm{h}j}\left[(m_{i}+m_{j})/M_{\mathrm{sub}}({<}r_{\mathrm{h}j})\right]^{1/3}$.
@@
where $N_{i}$ denotes the number and $m_{i}$ the mass of high-mass stars. Analogously, we denote by $N_{j}$, $m_{j}$, $r_{\mathrm{h}j}$ and $\langle v^{2}_{j}\rangle^{1/2}$ the number, mass, (3D) half-light radius and (3D) velocity dispersion of the field population. For massive–low mass binaries, the field population is the population of low-mass stars. For massive–massive binaries, the field population coincides with the population of massive stars, and we set $N_{j}=N_{i}-1$. The radius $a_{\mathrm{max}}$ here is chosen to be the largest semi-major axis for which the binary identification criterion employed in the simulations holds, assuming a binary located at the half-light radius $r_{\mathrm{h}j}$.
From Eq. 9, we see that the fraction of dynamically formed binaries scales with the (proxy) phase space density of the field population, $\propto N_{j}/(r_{\mathrm{h}j}^{3}\langle v^{2}_{j}\rangle^{3/2})$. The latter increases for the case of massive–massive binaries as the population of high-mass stars contracts and cools; hence, $f_{\mathrm{bin}}$ grows. Vice-versa, the (proxy) phase space density of low-mass stars decreases as the population expands and heats up, and $f_{\mathrm{bin}}$ drops. Dashed curves in Fig. 4 show the evolution of $f_{\mathrm{bin}}$ predicted by Eq. 9, in good agreement with the simulation results.
Figure 4: As the population of high-mass stars contracts and cools, the number of dynamically-formed binaries consisting of two massive stars increases. At the same time, as the population of low-mass stars expands and heats up, the number of dynamically-formed binaries consisting of a massive and a low-mass star decreases. Shown is the fraction $f_{\mathrm{bin}}$ of high-mass stars that are in binary systems relative to the total number of high-mass stars. Simulation results are shown as shaded bands, computed from a sample of $2000$ $N$-body realizations of the UMa3/U1 model. Dashed curves show the model predictions of Eq. 9.
@@ 4.3 Sensitivity to the dynamical-to-stellar mass ratio
The models described in the previous section assumed an initial dynamical-to-stellar mass ratio of $\Upsilon_{\mathrm{dyn}}=10$. However, Eq. 5 shows that the relaxation time, and hence the effects of collisional processes on the system, depend on the value of $\Upsilon_{\mathrm{dyn}}$. For $\Upsilon_{\mathrm{dyn}}\rightarrow\infty$, the system becomes collisionless, whereas for $\Upsilon_{\mathrm{dyn}}\rightarrow 1$, the dynamics are those of a classical star cluster. To study this dependence on $\Upsilon_{\mathrm{dyn}}$, we run a series of simulations varying the initial dynamical-to-stellar mass ratio over a range of $0.5\leq\log_{10}\Upsilon_{\mathrm{dyn}}\leq 2.5$. Note that we only study models that are dark matter-dominated, as our $N$-body setup does not capture the dynamical effects of stars on the dark matter cusp which would become more relevant as $\Upsilon_{\mathrm{dyn}}$ decreases (see, e.g., Zhang & Amaro Seoane 2025).
In Figure 5, we show the ratio between the half-light radii of the low-mass and high-mass stellar populations after $10\,\mathrm{Gyr}$ of evolution, for different values of the initial dynamical-to-stellar mass ratio $\Upsilon_{\mathrm{dyn}}$. As expected, for large values of $\Upsilon_{\mathrm{dyn}}$, there is no appreciable difference between the half-light radii of the two stellar populations: the system is collisionless. At the other extreme, for $\Upsilon_{\mathrm{dyn}}\sim 3$ and $10\,\mathrm{Gyr}$ of evolution, the stellar half-light radius of the low-mass population is three times, 2 times and 50 per cent larger than that of the population of low-mass stars for the case of the UMa3/U1, Del1 and Eri3 model, respectively. Error bars span the $16^{\mathrm{th}}$ to $84^{\mathrm{th}}$ percentile of the underlying distribution of $N$-body relizations. As before, Poisson noise renders any detection of mass segregation highly challenging for systems with a low number of member stars such as UMa3/U1.
Figure 5: Stellar mass segregation plays a substantial role for the dynamical evolution of systems akin to UMa3/U1 if their dynamical-to-stellar mass ratios $\Upsilon_{\mathrm{dyn}}$ are smaller than ${\sim}50$. Shown is the median ratio of half-light radii between the populations of low-mass and high-mass stars after $10\,\mathrm{Gyr}$ of evolution for different (initial) dynamical-to-stellar mass ratios $\Upsilon_{\mathrm{dyn}}$, computed over all $N$-body realizations (see text). Error bars span the $16^{\mathrm{th}}$ to $84^{\mathrm{th}}$ percentile of the underlying distribution.

### v2

@@ 3.2 Initial conditions
$\rho_{\star}(r)=\left[M_{\star}/(8\pi r_{\star}^{3})\right]~\exp(-r/r_{\star})~,$ (7)
@@
$\rho_{\mathrm{sub}}(r)=\left[M_{\mathrm{sub}}/(2\pi r_{\mathrm{sub}}^{3})\right]~(r/r_{\mathrm{sub}})^{-1}(1+r/r_{\mathrm{sub}})^{-3}~.$ (8)
@@
NN-body realizations. We generate equilibrium $N$-body realizations of Eq. 7 in the combined potential of the dark matter and stellar components using using the Eddington-inversion code nbopy (Errani & Peñarrubia, 2020), available online [^6] https://github.com/rerrani/nbopy. We assume that both stellar components have isotropic velocities in the initial conditions. The dark matter subhalo (Eq. 8) is modeled as an analytical background potential. To reduce the impact of Poisson noise on our analysis, for each choice of dynamical-to-stellar mass ratio $\Upsilon_{\mathrm{dyn}}$, we generate $200$ realizations of our UMa3/U1 model, $40$ for Del1, and $10$ for Eri3.
@@ 3.3 $N$-body code
For the UMa3/U1 models, we update the particle tree responsible for the long-range forces with a step size of $\Delta t_{\mathrm{L}}=2^{-10}r_{\mathrm{sub}}^{3/2}(GM_{\mathrm{sub}})^{-1/2}$. With this choice, the period of a circular orbit at the initial half-light radius $r_{\mathrm{h}}/r_{\mathrm{sub}}=1/500$ of a particle that is only subject to long-range forces (including the force due to the analytical dark matter potential) is resolved by ${\approx}270$ steps for the models with $\Upsilon_{\mathrm{dyn}}=10$. To test for numerical convergence, we have decreased and increased this value by factors of 4, with no qualitative impact on our results. Following the recommendation of Wang et al. (2020a, see their Eq. 12 and 41), we choose as reference radii for the separation of short- and long-range forces the values $r_{\mathrm{in,ref}}=\Delta t_{\mathrm{L}}\sigma_{1D}$ and $r_{\mathrm{out,ref}}=10\,r_{\mathrm{in,ref}}$, where by $\sigma_{1D}$ we denote the $N$-body system’s velocity dispersion. Slow-down regularization through sdar is applied to particles below a radius $r_{\mathrm{bin}}=0.8\,r_{\mathrm{in,ref}}$ (the default value in petar). To test for convergence, we have decreased this radius by a factor of 8, again with no qualitative impact on our results.
For the Del1 models, we use the same tree time step and reference radii as for UMa3/U1. For Eri3, we instead use a shorter tree time step of $\Delta t_{\mathrm{L}}=2^{-12}r_{\mathrm{sub}}^{3/2}(GM_{\mathrm{sub}})^{-1/2}$, which results in better performance at that $N$-body particle number by reducing the number of particles within $r_{\mathrm{in,ref}}$.
@@
4 Simulation results
We can now turn our attention to the results of our $N$-body experiments. In Sec. 4.1, 4.2 and 4.3, we discuss simulations of UMa3/U1, Del1 and Eri3 where we embed the stellar component in a static dark matter halo. Then, in Sec. 4.4, we model the effect of tides by allowing the underlying dark matter potential to evolve with time.
@@ 4.1 Simulations in a static dark matter halo
Fig. 2 shows simulation snapshots of the UMa3/U1 (top), Del1 (center) and Eri3 (bottom) models, evolved for $10\,\mathrm{Gyr}$ in a static dark matter halo. These models have an initial dynamical-to-stellar mass ratio of $\Upsilon_{\mathrm{dyn}}=10$ within the (initial) half-light radius. Individual low-mass (high-mass) stars are shown as light blue (grey) points, respectively. As time progresses, the population of low-mass stars expands, while the population of high-mass stars contracts: Even though the system is highly dark matter-dominated, the stellar population undergoes mass segregation [^7] Mass segregation occurs as close encounters between stars provide a means for the exchange of energy. The system’s tendency toward equipartition of energy results in low-mass stars preferentially moving to less-bound orbits (the low-mass population expands), whereas the high-mass stars move toward more tightly bound orbits (the high-mass population contracts), see, e.g., Spitzer (1987, chapter 1.3). driven by collisional relaxation. Blue and black circles in Fig. 2 show the median half-light radii of the low-mass (high-mass) population, with the median computed from the sample of all $N$-body realizations. Mass segregation progresses fastest for UMa3/U1 and slowest for Eri3, consistent with the relaxation times estimated in Fig. 1.
The detailed time evolution of the models with $\Upsilon_{\mathrm{dyn}}=10$ is shown in Fig. 3. The top panels show the evolution of the (median) half-light radii of the populations of low-mass (blue) and high-mass (black) stars. For the case of UMa3/U1, after $10\,\mathrm{Gyr}$ of evolution, the half-light radius of the population of low-mass stars is twice as large as the half-light radius of high-mass stars. For Del1, the difference reduces to ${\sim}40$ per cent, and for Eri3 to ${\sim}10$ per cent. The evolution of individual $N$-body realizations is shown in lighter shades in the background: For systems akin to UMa3/U1, Poisson noise driven by the low number of stars in the system substantially complicates a detection of mass segregation. The bottom panels of Fig. 3 show the evolution of the (3D) stellar velocity dispersion $\langle v^{2}\rangle=\sum v^{2}/N_{\star}$. As the half-light radius of the population of low-mass stars expands, the velocity dispersion heats up. Vice versa, as the population of high-mass stars contracts, the population cools down. This behaviour is expected for gravitationally subdominant stellar components that are deeply embedded within a dark matter subhalo (see footnote 10 for a calculation of the velocity dispersion predicted by the virial theorem, in good agreement with the simulation results, and Peñarrubia et al. 2025 for a detailed discussion of gravothermal expansion). Note that this behaviour differs from that of fully self-gravitating stellar systems devoid of dark matter, which typically heat up as they contract (see, e.g., Lynden-Bell & Wood, 1968; Spitzer & Thuan, 1972).
@@ 4.2 Dynamical formation and disruption of binaries
As the population of high-mass stars contracts and cools, we note the dynamical formation and disruption of stellar binary systems in our simulations. We identify binaries as pairs of stars with Keplerian binding energy $E_{\mathrm{bin}}<0$ and an orbital period $T_{\mathrm{bin}}$ that is shorter than the (circular) period $T_{\mathrm{com}}$ of the pair’s center of mass within the dark matter subhalo. Expressed in terms of densities, this condition implies that the mean stellar density within the semimajor axis $a_{\mathrm{bin}}$ of a binary exceeds the mean density of the dark halo within a radius equal to that of the binary’s center of mass , i.e $(m_{i}+m_{j})/a_{\mathrm{bin}}^{3}>M_{\mathrm{sub}}({<}r_{\mathrm{com}})/r_{\mathrm{com}}^{3}$, where by $m_{i}$ and $m_{j}$ we denote the masses of the two binary components. We count the number of binaries at each snapshot in the simulation. For better statistics, we stack $2000$ realizations of the UMa3/U1 model with an initial dynamical-to-stellar mass ratio of $\Upsilon_{\mathrm{dyn}}=10$. Note that our initial conditions are created by drawing individual stars from the underlying distribution function. Any binaries present in the initial conditions just arise from this random sampling, resulting in a binary fraction that is substantially lower than current observational constraints [^8] Spectroscopic studies suggest that the fraction of stellar binaries in dwarf galaxies is substantial: Spencer et al. (2017) infer a binary fraction of ${\sim}50$ per cent for Draco and ${\sim}80$ per cent in Ursa Minor.. We adopt this modeling choice for the sake of simplicity.
@@
Fig. 4 shows the binary fraction $f_{\mathrm{bin}}\equiv N_{\mathrm{bin},i}/N_{i}$, defined here as the number $N_{\mathrm{bin},i}$ of high-mass stars that are in binary systems, normalized by the total number of high-mass stars $N_{i}$. A grey band shows the binary fraction considering only pairs of two high-mass stars, whereas the blue band corresponds to binaries that consist of a high-mass star and a low-mass star. As the population of high-mass stars contracts and cools due to mass segregation, the number of dynamically formed massive–massive binaries increases. At the same time, as the population of low-mass stars expands and heats up, the number of massive–low mass binaries drops. The stellar binaries in our simulations dynamically form and later dissolve again, consistent with the results of Petit & Henon (1986) who show that all gravitational captures (in three-body systems) are temporary configurations (see also Peñarrubia 2023; Peñarrubia et al. 2024).
@@
Some intuition in the processes driving the formation and disruption of dynamical binaries can be gained from the statistical theory of gravitational capture, developed to estimate the number of gravitationally trapped (massless) tracer particles around a point-mass perturber orbiting in a smooth dark matter halo (Peñarrubia, 2023). In our collisional $N$-body simulations, all stars are massive particles, and complicated three- or many-body interaction between stars and the halo likely contribute to the formation and disruption processes. Nevertheless, as we will show in the following, the statistical theory of gravitational capture provides accurate estimates for the binary fractions found in our simulations. Building upon Eq. 4 in Peñarrubia (2021), we estimate the binary fraction through [^9] We obtain our Eq. 9 from Eq. 4 of Peñarrubia (2021) by approximating the mean number density of field stars trough $\bar{n}_{j}=3N_{j}/(8\pi r_{\mathrm{h}j}^{3})$. We then substitute the mass of the perturber by the mass of the binary system, $m_{i}+m_{j}$. Finally, we compute the number of bound field stars within a radius $a_{\mathrm{max}}=r_{\mathrm{h}j}\left[(m_{i}+m_{j})/M_{\mathrm{sub}}({<}r_{\mathrm{h}j})\right]^{1/3}$.
@@
where $N_{i}$ denotes the number and $m_{i}$ the mass of high-mass stars. Analogously, we denote by $N_{j}$, $m_{j}$, $r_{\mathrm{h}j}$ and $\langle v^{2}_{j}\rangle^{1/2}$ the number, mass, (3D) half-light radius and (3D) velocity dispersion of the field population. For massive–low mass binaries, the field population is the population of low-mass stars. For massive–massive binaries, the field population coincides with the population of massive stars, and we set $N_{j}=N_{i}-1$. The radius $a_{\mathrm{max}}$ here is chosen to be the largest semimajor axis for which the binary identification criterion employed in the simulations holds, assuming a binary located at the half-light radius $r_{\mathrm{h}j}$.
From Eq. 9, we see that the fraction of dynamically formed binaries scales with the (proxy) phase space density of the field population, $\propto N_{j}/(r_{\mathrm{h}j}^{3}\langle v^{2}_{j}\rangle^{3/2})$. The latter increases for the case of massive–massive binaries as the population of high-mass stars contracts and cools; hence, $f_{\mathrm{bin}}$ grows. Vice versa, the (proxy) phase space density of low-mass stars decreases as the population expands and heats up, and $f_{\mathrm{bin}}$ drops. Dashed curves in Fig. 4 show the evolution of $f_{\mathrm{bin}}$ predicted by Eq. 9, which is in good agreement with the abundance of binaries identified in the $N$-body simulations.
In our simplified setup, most dynamically formed binaries are wide and thus contribute minimally to the systemic velocity dispersion. As a result, they may evade detection in spectroscopic surveys but can imprint a distinct signal in the stellar two-point correlation function (Longhitano & Binggeli, 2010; Kervick et al., 2022; Safarzadeh et al., 2022). Given that the statistics of wide binaries have been proposed as a sensitive probe of dark matter substructure (Peñarrubia et al., 2016; Shariat et al., 2025), our simulation results, suggesting the dynamical formation and disruption of (wide) binaries in dark matter-dominated dwarf galaxies, emphasize the importance of detailed modeling of the dynamical processes that govern the abundance of wide binaries in these systems.
Figure 4: As the population of high-mass stars contracts and cools, the number of dynamically formed binaries consisting of two massive stars increases. At the same time, as the population of low-mass stars expands and heats up, the number of dynamically formed binaries consisting of a massive and a low-mass star decreases. Shown is the fraction $f_{\mathrm{bin}}$ of high-mass stars that are in binary systems relative to the total number of high-mass stars. Simulation results are shown as shaded bands, computed from a sample of $2000$ NN-body realizations of the UMa3/U1 model. Dashed curves show the model predictions of Eq. 9.
@@ 4.3 Sensitivity to the dynamical-to-stellar mass ratio
The models described in the previous sections assumed an initial dynamical-to-stellar mass ratio of $\Upsilon_{\mathrm{dyn}}=10$. However, Eq. 5 shows that the relaxation time, and hence the effects of collisional processes on the system, depend on the value of $\Upsilon_{\mathrm{dyn}}$. For $\Upsilon_{\mathrm{dyn}}\rightarrow\infty$, the system becomes collisionless, whereas for $\Upsilon_{\mathrm{dyn}}\rightarrow 1$, the dynamics are those of a classical star cluster. To study this dependence on $\Upsilon_{\mathrm{dyn}}$, we run a series of simulations, varying the initial dynamical-to-stellar mass ratio over the range $0.5\leq\log_{10}\Upsilon_{\mathrm{dyn}}\leq 2.5$. Note that we only study models that are dark matter-dominated, as our $N$-body setup does not capture the dynamical effects of stars on the dark matter cusp, which would become more relevant as $\Upsilon_{\mathrm{dyn}}$ decreases (see, e.g., Zhang & Amaro Seoane 2025).
In Fig. 5, we show the ratio between the half-light radii of the low-mass and high-mass stellar populations after $10\,\mathrm{Gyr}$ of evolution, for different values of the initial dynamical-to-stellar mass ratio $\Upsilon_{\mathrm{dyn}}$. As expected, for large values of $\Upsilon_{\mathrm{dyn}}$, there is no appreciable difference between the half-light radii of the two stellar populations: the system is collisionless. At the other extreme, for $\Upsilon_{\mathrm{dyn}}\sim 3$ and $10\,\mathrm{Gyr}$ of evolution, the stellar half-light radius of the low-mass population is three times, 2 times and 50 per cent larger than that of the population of low-mass stars for the cases of the UMa3/U1, Del1 and Eri3 models, respectively. Error bars span the $16^{\mathrm{th}}$ to $84^{\mathrm{th}}$ percentiles of the underlying distribution of $N$-body relizations. As before, Poisson noise renders any detection of mass segregation highly challenging for systems with a low number of member stars such as UMa3/U1.
Figure 5: Stellar mass segregation plays a substantial role in the dynamical evolution of systems akin to UMa3/U1 if their dynamical-to-stellar mass ratios $\Upsilon_{\mathrm{dyn}}$ are smaller than ${\sim}50$. Shown is the median ratio of half-light radii between the populations of low-mass and high-mass stars after $10\,\mathrm{Gyr}$ of evolution for different (initial) dynamical-to-stellar mass ratios $\Upsilon_{\mathrm{dyn}}$, computed over all $N$-body realizations (see text). Error bars span the $16^{\mathrm{th}}$ to $84^{\mathrm{th}}$ percentiles of the underlying distribution.